|
| |||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||
| Stack Data | ||||||||||||
|---|---|---|---|---|---|---|---|---|---|---|---|---|
| Start | Stop | ee50 | ee80 | Hefi Focus | Mars Focus | Outside Temp | Dome Temp | Truss Temp | Wind Direction | Wind Speed | Stack Type | |
| 05:10 | 08:37 | 0.48 | 0.75 | -520 | +7690 | 15.8 | 13.5 | 265 | 7 | First of Night | ||
| Problem Reports | ||||
|---|---|---|---|---|
| PR Number | Times | Time Lost | Description | Response |
| 3944 | 03:00-04:00 | 01:00 | This is not currently a problem, but it's filed as a reminder to anyone who goes in the CCAS tower to make sure that the fiber has not been pulled from the AOA/MARS light source when any work up there is done. | |
| 1703 | 05:00-05:00 | 00:00 | Removed: 8-1, 21-4, 85-4, 79-5, 79-1, 70-1, 70-2, 70-6 | |
| Ops Comments | none |
|---|
| TO Comments | John Good took measurements at CCAS to center the structure at the tower. Matt fixed light source & mirror locations file so that AOA routine works once again. Was able to get a great stack tonight. Found all mirrors. |
|---|
| RA Comments | none |
|---|
| Observations | |||||||||||||||
|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|
| File | Target/QID | Exp Start | Exposure | Overhead | Program | Instrument Configuration | Data Type | IQ | Dimm | SkySBrt | Skytrans | SNR | Comments | Charge | |